# Metric Distortion in Peer Selection

### v1

@@ Abstract
In the metric distortion problem, a set of voters and candidates lie in a common metric space, and a committee of $k$ candidates is to be elected. The goal is to select a committee with a small social cost, defined as an increasing function of the distances between voters and selected candidates, but a voting rule only has access to voters’ ordinal preferences. The distortion of a rule is then defined as the worst-case ratio between the social cost of the selected set and the optimal set, over all possible preferences and consistent distances.
We initiate the study of metric distortion when voters and candidates coincide, which arises naturally in peer selection, and provide tight results for various social cost functions on the line metric. We consider both utilitarian and egalitarian social cost, given by the sum and maximum of the individual social costs, respectively. For utilitarian social cost, we show that the simple voting rule that selects the $k$ middle agents achieves a distortion that varies between $1$ and $2$ as $k$ varies between $1$ and $n$ when the cost of an individual is the sum of their distances to all selected candidates (additive aggregation). When the cost of an individual is their distance to their $q$th closest candidate ($q$-cost), we provide positive results for $q=k=2$ but mostly show that negative results for general elections carry over to our restricted setting: No constant distortion is possible when $q\leq k/2$ and no distortion better than $3/2$ is possible for $q\geq k/2+1$. For egalitarian social cost, a rule that selects extreme agents achieves the best-possible distortion of $2$ for additive cost and $q$-cost with $q>k/3$, whereas no constant distortion is possible for $q\leq k/3$. Our results suggest that having a common set of voters and candidates allows for better constants compared to the general setting, but cases in which no constant is possible in general remain hard under this restriction.
@@ 1 Introduction
A fundamental problem in social choice is the aggregation of individual preferences, expressed as rankings over a set of candidates, into a social preference consisting of a subset of elected candidates. For centuries, social choice theorists have proposed several axioms to capture desirable properties that these aggregation or voting rules should guarantee, usually leading to strong impossibility results (Arrow, 1963; de Condorcet, 1785; Gibbard, 1973; Satterthwaite, 1975).
@@
As an alternative approach, attempting to quantify the extent to which a certain voting rule is able to faithfully translate the voter preferences into the selected committee, Procaccia and Rosenschein (2006) introduced the notion of distortion of a rule. The underlying assumption is that a voter’s (dis)affinity with a candidate can be represented by a certain cost, and voters’ rankings are the expression of these cardinal preferences. The cost of a committee for a voter is then defined by aggregating the costs of the committee members, and the overall social cost of the committee by aggregating the costs for all voters. The distortion then corresponds to the worst-case ratio between the social cost of the selected committee and that of the optimal committee, over all possible preferences and consistent costs.
@@
The study of the distortion of voting rules has usually focused on two ways of modeling the social cost: utilitarian and egalitarian (Caragiannis and Procaccia, 2011; Goel et al., 2018; Caragiannis et al., 2017). In the utilitarian case, the social cost is defined as the sum of the individual costs of the voters, ensuring that all voters’ costs contribute equally to the objective. In contrast, the egalitarian social cost considers the maximum individual cost among all voters, aiming to capture a notion of fairness where no voter is excessively disadvantaged.
@@
In voting theory, it is common to assume that voters’ preferences are not fully arbitrary but enjoy some structural properties. A relevant line of work has indeed sought structural restrictions that are natural and have powerful implications, such as single-peaked (Black, 1948) or single-crossing (Mirrlees, 1971); see Elkind et al. (2022) for a survey. A rather general framework among these is that of spatial or metric voting, where voters and candidates are assumed to lie in a common low-dimensional metric space and voters’ costs correspond to their distance to each candidate (Aziz, 2020; Jessee, 2012; Enelow and Hinich, 1984; Merrill and Grofman, 1999). For instance, a line metric is commonly employed to capture political affinity on the left-right spectrum, whereas geographical distances are represented in a two-dimensional space.
@@
This structural assumption naturally fits in the metric distortion framework: The distances to candidates fully define the social cost of a committee, but the voting rules only receive their expression as preference rankings. Since preferences are restricted in this model, improved bounds on the distortion of voting rules have been established. Notably, a tight distortion bound of $3$ has been established for any single-winner deterministic voting rule (Anshelevich et al., 2018; Kizilkaya and Kempe, 2022; Gkatzelis et al., 2020). Extending distortion to multi-winner elections requires defining how a voter’s cost is aggregated over the selected committee. Two ways have been considered in the literature: the additive cost, where a voter’s cost is the sum of their distances to all members of the committee (Babashah et al., 2024), and the $q$-cost, where the cost is determined by their distance to their $q$-th closest committee member (Caragiannis et al., 2022b; Chen et al., 2020).
@@
Work on metric distortion has so far focused on the case where voters and candidates constitute disjoint sets, which constitutes a natural model for large-scale elections. However, in many decision-making scenarios, a group of agents aims to elect a subset of their own members. One can think, for example, of a political organization selecting a committee. Each member ranks others according to their political affinity and the organization aims to select a committee that represents the variety of preferences of its members. Since the voting rule only receives ordinal preferences, a small distortion constitutes a suitable objective to ensure a close-to-optimal outcome under this limited information. In general, this situation arises in the context of peer selection, where individuals evaluate each other to choose a group for governance, leadership, or resource allocation. Further examples include academic hiring and promotions, student representative elections, self-organized committees in cooperatives, and local governance selection.
@@
While peer selection rules have been extensively studied in other contexts, particularly in terms of the effect of strategic behavior (e.g. Holzman and Moulin, 2013; Alon et al., 2011; Caragiannis et al., 2022a), little is known regarding their ability to accurately reflect agents’ ordinal preferences. On the other hand, previous work on metric distortion for single-selection has often parameterized an election via its decisiveness, corresponding to the maximum ratio between a voter’s distance to their top choice and to any other candidate (Anshelevich and Postl, 2017; Gkatzelis et al., 2020). These works have motivated this parameter with the fact that it becomes zero in the peer selection setting as each agent becomes their own top choice.
However, directly considering a common set of voters and candidates constitutes a structural modification to the problem that has not been considered so far.
@@ 1.1 Our Contributions and Techniques
We initiate the study of metric distortion when the set of voters and candidates coincide and bound the distortion achievable by voting rules selecting $k$ out of $n$ agents on the line metric for several social costs; see Table 1 for a summary of our results.
additive $q$-cost $q\leq\frac{k}{3}$ $\frac{k}{3}<q\leq\frac{k}{2}$ $\frac{k}{2}<q\leq k$ $q=k=2$ utilitarian $1+\sqrt{1+\frac{2}{k}}$; $\frac{7}{3}+\frac{4}{k}\big{(}\sqrt{2}-\frac{4}{3}\big{)}$ [BKSS]   $\infty$ [CSV]   $3$; $3$ [CSV] $1$; $\frac{2}{k}\Big{(}n-\sqrt{2n\big{\lfloor}\frac{n-k}{2}\big{\rfloor}}\Big{)}$ [T. 1]   $\infty$ [T. 2  $2-\frac{k-q}{4q-k-3}^{(*)}$; $3$ [T. 3, CSV] $2$; $2$ ($n$ even) [P. 1] $\frac{4}{3}$; $\frac{4}{3}$ ($n$ odd) [T. 4] egalitarian $\infty$ [CSV]   $3$; $3$ [CSV]   $\frac{3}{2}-\frac{1}{k}$; $\frac{3}{2}-\frac{1}{2(k-1)}$ ($k$ even) [T. 5] $\infty$ [T. 6]   $2$; $2$ [T. 7  $\frac{3}{2}-\frac{1}{k}$; $\frac{3}{2}-\frac{1}{k(k-1)}$ ($k$ odd) [T. 5]
Table 1: Our and previous bounds on the distortion that voting rules can achieve in different settings. Values before and after the semicolon represent lower and upper bounds for the corresponding setting, respectively. Lower bounds take worst-case number of agents $n$. When the lower bound can be made arbitrarily large, we just write $\infty$ for simplicity. The number in square brackets refers to the theorem (T.) or proposition (P.) where this bound is shown; the letters in square brackets refer to the paper where a bound is taken from: BKSS is Babashah et al. (2024) and CSV is Caragiannis et al. (2022b). In particular, bounds in gray correspond to the previously studied setting with disjoint voters and candidates for comparison, either under a general metric [CSV] or under the line metric [BKSS]. The upper bound for utilitarian $q$-cost marked with ()(*)( ∗ ) is only valid when $q\geq\frac{k}{2}+1$, which is slightly stronger than the general condition $q>\frac{k}{2}$ on that column.
We start by observing a simple yet strong property of metric voting on the line with a single set of voters and candidates that follows from previous work (Elkind and Faliszewski, 2014; Babashah et al., 2024): We can fully compute the order of the agents from their rankings. This constitutes a powerful tool for the design of our mechanisms, as in the following we can always take this order as given.
@@
We first consider the utilitarian social cost, in which the social cost of a committee is defined as the sum of all individual costs. Intuitively, selecting $k$ consecutive agents results in lower utilitarian social cost. In Section 3.1, we focus on the case of additive aggregation: The cost of a committee for a voter is given by the sum of all distances from the candidates to this voter. As a natural extension of the optimal rules for one or two agents, which select the median and closest-to-median agents, we consider a rule called Median Alternation that selects $k$ middle agents. We show that Median Alternation provides a distortion of at most $\frac{2}{k}\Big{(}n-\sqrt{2n\big{\lfloor}\frac{n-k}{2}\big{\rfloor}}\Big{)}$, which is close to $1$ when $k$ is small compared to $n$ and approaches $2$ as $k$ goes to $n$. Despite its simplicity, the analysis of this rule holds significant challenge. In short, we reduce any metric to another with only two locations and show the distortion for this class of metrics. We show that this reduction is possible without improving the distortion by establishing the existence of a non-improving direction of movement for each agent.
@@
In Section 3.2, we consider utilitarian $q$-cost, where the cost of a committee for an agent is given by the agent’s distance to their $q$th closest candidate in the committee. In Theorem 2, we show that no voting rule can provide a constant distortion when $q\leq\frac{k}{2}$, implying that this known impossibility from the setting with disjoint voters and candidates and a general metric space (Caragiannis et al., 2022b) remains in place in our restricted setting. To prove this bound, we partition all but $q$ agents into $\big{\lfloor}\frac{k}{q}\big{\rfloor}\geq 2$ sets and consider two metrics that differ in the position of the remaining $q$ agents: relatively close to the other agents in one metric; very far in the other. Intuitively, selecting these $q$ agents leads to an unbounded distortion in the former case but is necessary for a bounded distortion in the latter. For $q>\frac{k}{2}$, the existence of rules with distortion $3$ follows from a general result by Caragiannis et al. (2022b). We provide a lower bound, that varies between $\frac{3}{2}$ and $2$ as $q$ varies between $\frac{k}{2}+1$ and $2$, by considering three different metrics consistent with the same rankings and showing that, in one of them, there are $q$ agents in one extreme that cannot be consistently selected. We finally take a closer look at the case with $k=q=2$, where a best-possible distortion of $2$ can be achieved by selecting the median agents when $k$ is even. For odd $k$, we show that a rule selecting a couple of agents—a pair of agents who prefer each other over all other agents—among the five middle agents achieves an improved distortion of $\frac{4}{3}$, which is again best-possible. The Favorite Couple rule leverages two key principles: (1) selecting agents close to the median so as to balance overall distances from agents on each side of the median, and (2) selecting consecutive agents with a small distance between them since this distance is part of the cost of all agents. This intuition of selecting consecutive agents that are as close to each other while also being close to the median in principle holds for larger $k$, but determining how tightly a group of $k$ agents is clustered based solely on ordinal rankings remains a challenge.
@@
In Section 4.3, we show that $k$-Extremes attains a distortion of $2$ for $q$-cost as long as $q>\frac{k}{3}$. To do so, we prove that the social cost of the set selected by this rule is at most the distance from the agent closest to the center to their nearest extreme, and bound the social cost of the optimal set from below by half of this distance. We provide a matching lower bound by revisiting the instance used for the additive case. Finally, we show that no constant distortion is possible when $q\leq\frac{k}{3}$, again implying that the general impossibility result of Caragiannis et al. (2022b) still holds in our setting. In the worst-case instances, we partition the agents into $\big{\lfloor}\frac{k}{q}\big{\rfloor}$ sets and consider two symmetric distance metrics where all but one set are placed at a unit distance from one another and two sets in one extreme are at the same location. We show that no rule can pick $q$ agents from each location.
@@ 1.2 Further Related Work
In the non-metric framework, Caragiannis and Procaccia (2011) showed that the distortion of any voting rule is at least $\Omega(m^{2})$ and that simple rules such as Plurality achieve a distortion of at most $O(m^{2})$, where $m$ is the number of candidates.
In the metric framework, Anshelevich et al. (2018) established a general lower bound of $3$ on the distortion of any deterministic voting rule. They also analyzed the distortion of common voting rules such as Majority, Borda, and Copeland, showing that the latter achieves the lowest distortion of $5$ among them. Goel et al. (2017) disproved a conjecture by Anshelevich et al. regarding a better-than-$5$ distortion of the Ranked Pairs rule and introduced the notion of fairness ratio of a rule, which captures the egalitarian social cost as a special case. These results were later improved by Munagala and Wang (2019), who extended the analysis to uncovered set rules and reduced the upper bound to $4.236$. Gkatzelis et al. (2020) closed the gap by improving this bound to $3$, and showed the validity of this bound in terms of fairness ratio and thus egalitarian social cost.
Randomized voting rules have also been extensively explored in the metric framework (Pulyassary and Swamy, 2021; Fain et al., 2019). The best-known upper bound for a randomized voting rule was recently obtained by Charikar et al. (2024), who showed that a carefully designed randomization over existing and novel voting rules achieves a distortion of at most $2.753$. As of lower bounds, Charikar and Ramakrishnan (2022) disproved a conjecture by Goel et al. (2017) regarding the existence of a randomized voting rule with distortion $2$, by constructing instances whose distortion approaches $2.113$ as the number of candidates grows.
@@
In the study of metric distortion for multi-winner voting, various objective functions have been proposed to capture the cost incurred by each voter for the elected committee (Elkind et al., 2017; Faliszewski et al., 2017). A foundational result by Goel et al. (2018) showed that, for the additive cost function, iterating a single-winner voting rule with distortion $\delta$ for $k$ rounds produces a $k$-winner committee with the same distortion.
Chen et al. (2020) studied the $1$-cost objective in the metric framework when each voter casts a vote for a single candidate. They proposed a deterministic rule with a tight distortion of $3$ and a randomized rule with a distortion of $3-\frac{2}{m}$. More generally, Caragiannis et al. (2022b) introduced the $q$-cost objective, where a voter’s cost for a committee is determined by the distance to their $q$-th closest member. They showed that the distortion is unbounded for $q\leq\frac{k}{3}$ and linear in $n$ for $\frac{k}{3}<q\leq\frac{k}{2}$. For $q>\frac{k}{2}$, they presented a non-polynomial voting rule that achieves a distortion of $3$ and a polynomial rule with a distortion of $9$. They discussed how these upper bounds for $q>\frac{k}{2}$ and the unbounded distortion for $q\leq\frac{k}{3}$ carry over to egalitarian social cost, but interestingly showed that a constant distortion is possible for this objective when $\frac{k}{3}<q\leq\frac{k}{2}$. Kizilkaya and Kempe (2022) later proposed a polynomial-time rule with a distortion of $3$. Recently, Babashah et al. (2024) studied the distortion of multi-winner elections with additive cost on the line, devising a rule with a distortion of roughly $\frac{7}{3}$.
Caragiannis et al. (2017) studied distortion in multi-winner voting for the non-metric framework, defining a voter’s utility for a committee as the highest utility derived from any of its members. They proposed a rule achieving a distortion of $1+\frac{m(m-k)}{k}$ for deterministic committee selection when selecting $k$ out of $m$ candidates.
@@
A specialized setting in metric voting considers single-peaked and $1$-Euclidean preferences, where both voters and alternatives are embedded on the real line (Black, 1948; Moulin, 1980; Miyagawa, 2001; Fotakis et al., 2016; Fotakis and Gourvès, 2022; Voudouris, 2023; Ghodsi et al., 2019). In particular, the work of Fotakis et al. (2024) investigated the distortion of deterministic algorithms for $k$-committee selection on the line under the $1$-cost objective, leveraging additional distance queries.

### v2

@@ Abstract
In the metric distortion problem, a set of voters and candidates lie in a common metric space, and a committee of $k$ candidates must be elected. The objective is to minimize a social cost, defined as a function of the distances between voters and their chosen representatives, while the voting rule only has access to ordinal preferences. The distortion of a rule is the worst-case ratio between the social cost of its outcome and that of the optimal committee, taken over all consistent preferences and metrics.
We initiate the study of metric distortion in peer selection, where voters and candidates coincide. We consider four objectives, obtained by combining two aggregation rules with two types of social cost. Under additive aggregation, an individual’s cost is the sum of their distances to all committee members; under $q$-cost, it is their distance to the $q$th closest member. The overall social cost is either utilitarian, given by the sum of all individual costs, or egalitarian, given by the maximum individual cost. Surprisingly, we find that even on the line metric, peer selection retains much of the hardness of the general case: Lower bounds remain strictly larger than one for all objectives, and cases where bounded distortion is impossible in general remain so here as well. On a positive note, cases with bounded distortion in the general setting achieve better constants in peer selection. For utilitarian cost, selecting the $k$ middle agents achieves a distortion between $1$ and $2$ under additive aggregation. Under $q$-cost, we show positive results for $q=k=2$, but impossibility results largely carry over. For egalitarian cost, selecting the extremes yields an optimal distortion of $2$ under additive aggregation and for $q$-cost with $q>k/3$, while no bounded distortion is possible when $q\leq k/3$. Overall, our results show that peer selection on the line metric admits better constants than the general case, yet fundamental hardness barriers persist.
@@ 1 Introduction
A fundamental problem in social choice is the aggregation of individual preferences, expressed as rankings over a set of candidates, into a social preference consisting of a subset of elected candidates. For centuries, social choice theorists have proposed several axioms to capture desirable properties that these aggregation or voting rules should satisfy, usually leading to strong impossibility results (Arrow, 1963; de Condorcet, 1785; Gibbard, 1973; Satterthwaite, 1975).
@@
As an alternative approach, attempting to quantify the extent to which a certain voting rule can faithfully translate voters’ preferences into the selected committee, Procaccia and Rosenschein (2006) introduced the notion of distortion of a rule. The underlying assumption is that a voter’s (dis)affinity with a candidate can be represented by a certain cost, and voters’ rankings express these cardinal preferences. The cost of a committee for a voter is then defined by aggregating the costs of the committee members, and the overall social cost of the committee by aggregating the costs for all voters. The distortion corresponds to the worst-case ratio between the social cost of the selected committee and that of the optimal committee, over all possible preferences and consistent metrics.
@@
The study of the distortion of voting rules has usually focused on two ways of modeling the social cost: utilitarian and egalitarian (Caragiannis and Procaccia, 2011; Goel et al., 2018; Caragiannis et al., 2017). In the utilitarian case, the social cost is defined as the sum of the individual costs of the voters, so that all voters’ costs contribute equally to the objective. In contrast, the egalitarian social cost considers the maximum individual cost among all voters, capturing a notion of fairness where no voter is excessively disadvantaged.
@@
In voting theory, it is common to assume that voters’ preferences are not fully arbitrary but exhibit some structural properties. A relevant line of work has sought structural restrictions that are natural and have powerful implications, such as single-peaked (Black, 1948) or single-crossing (Mirrlees, 1971); see Elkind et al. (2022) for a survey. A rather general framework among these is that of spatial or metric voting, where voters and candidates are assumed to lie in a common low-dimensional metric space and voters’ costs correspond to their distance to each candidate (Aziz, 2020; Jessee, 2012; Enelow and Hinich, 1984; Merrill and Grofman, 1999). For instance, a line metric is commonly employed to capture political affinity on the left-right spectrum, whereas geographical preferences are typically modeled in two-dimensional spaces.
@@
This structural assumption naturally fits in the metric distortion framework: the distances to candidates fully define the social cost of a committee, but the voting rules only receive their expression as preference rankings. With this structural restriction, a tight distortion bound of $3$ has been shown for any single-winner deterministic voting rule (Anshelevich et al., 2018; Kizilkaya and Kempe, 2022; Gkatzelis et al., 2020). Extending distortion to multi-winner elections requires defining how a voter’s cost is aggregated over the selected committee. Two ways have been considered in the literature: the additive cost, where a voter’s cost is the sum of their distances to all members of the committee (Babashah et al., ), and the $q$-cost, where the cost is determined by their distance to their $q$th closest committee member (Caragiannis et al., 2022b; Chen et al., 2020).
@@
Work on metric distortion has so far focused on the case where voters and candidates constitute disjoint sets, which forms a natural model for large-scale elections. However, in many decision-making scenarios, a group of agents seeks to elect a subset of their own members. One can think, for example, of a political organization selecting a committee. Each member ranks others according to their political affinity, and the organization aims to select a committee that represents the variety of preferences of its members. Since the voting rule only receives ordinal preferences, a small distortion constitutes a suitable objective to ensure a close-to-optimal outcome under this limited information. In general, this situation arises in the context of peer selection, where individuals evaluate each other to choose a group for governance, leadership, or resource allocation. Further examples include academic hiring and promotion, student representative elections, self-organized committees in cooperatives, and local governance selection.
@@
While peer selection rules have been extensively studied in other contexts, particularly regarding the effect of strategic behavior (e.g. Holzman and Moulin, 2013; Alon et al., 2011; Caragiannis et al., 2022a), little is known about their ability to accurately reflect agents’ cardinal preferences. On the other hand, previous work on metric distortion for single-selection has often parameterized an election via its decisiveness, corresponding to the maximum ratio between a voter’s distance to their top choice and to any other candidate (Anshelevich and Postl, 2017; Gkatzelis et al., 2020). These works have motivated this parameter by the fact that it becomes zero in the peer selection setting, as each agent becomes their own top choice. However, directly modeling a common set of voters and candidates constitutes a structural change to the problem that has not been explored so far.
@@ 1.1 Our Contributions and Techniques
We initiate the study of metric distortion when the set of voters and candidates coincide and bound the distortion achievable by voting rules selecting $k$ out of $n$ agents on the line metric for several social costs. A summary of our results appears in Table 1.
additive $q$-costaaaaaaaaa $q\leq\frac{k}{2}$ $\frac{k}{2}<q\leq k$ utilitarian $1+\sqrt{1+\frac{2}{k}}$; $\frac{7}{3}+\frac{4}{k}\big(\sqrt{2}-\frac{4}{3}\big)$ [BKSS] $\infty$ [CSV] $3$; 33 [CSV] $1.0914$; $\tfrac{2}{k}\big(n-2\sqrt{n(n-k)}\big)^{(*)}$ [T. 1] $\infty$ [T. 3] $2-\tfrac{k-q}{4q-k-3}^{(**)}$; $3$ [T. 4, CSV] additive $q$-costaaaaaaaaa $q\leq\frac{k}{3}$ $\frac{k}{3}<q\leq k$ egalitarian $\cdot$; $2$ [BKSS] $\infty$ [CSV] $3$; 33 [CSV] $\tfrac{3}{2}-\tfrac{1}{k}$; $\tfrac{3}{2}-\tfrac{1}{2(k-1)}^{(*)}$ [T. 6] $\infty$ [T. 7] $2$; 22 [T. 8]
Table 1: Summary of our and previous bounds on the distortion that voting rules can achieve in different settings. Values before and after the semicolon represent lower and upper bounds, respectively. Gray entries correspond to the previously studied setting with disjoint voters and candidates, either under a general metric [CSV] or under the line metric [BKSS]. The upper bounds for utilitarian additive and egalitarian additive social costs marked with ()(*) correspond to the cases of even nkn-k and even $k$, respectively; the slightly different bounds for the odd cases are stated in the corresponding theorems. The upper bound for utilitarian $q$-cost marked with ()(**) in the text is valid for $q\geq\frac{k}{2}+1$ (slightly stronger than $q>\frac{k}{2}$). BKSS: Babashah et al. ; CSV: Caragiannis et al. (2022b).
We start by observing a simple yet powerful property of metric voting on the line with a single set of voters and candidates, which follows from previous work (Elkind and Faliszewski, 2014; Babashah et al., ): we can fully compute the order of the agents from their rankings. This constitutes a powerful tool for the design of our mechanisms, as in the following we can always take this order as given.
@@
We first consider the utilitarian social cost, in which the social cost of a committee is defined as the sum of all individual costs. Intuitively, selecting $k$ consecutive agents results in lower utilitarian social cost. In Section 3.1, we focus on the case of additive aggregation: the cost of a committee for a voter is given by the sum of all distances from the candidates to this voter. As a natural extension of the optimal rules for one or two agents, which select the median and closest-to-median agents, we consider a rule called Median Alternation that selects $k$ middle agents. We show that Median Alternation provides a distortion close to $1$ when $k$ is small compared to $n$ and approaches $2$ as $k$ goes to $n$. Despite its simplicity, the analysis of this rule poses significant challenges. In short, we reduce any metric to another with only two locations by showing the existence of a non-improving direction of movement for each agent, and then compute the worst-case distortion for this class.
We complement the upper bound described above with a lower bound of $1.0914$ for any $k\geq 3$. This implies that, even in this seemingly simple case, we cannot always select the optimal committee.
@@
In Section 3.2, we consider utilitarian $q$-cost, where the cost of a committee for an agent is given by the agent’s distance to their $q$th closest candidate in the committee. In Theorem 3, we show that no voting rule can provide a constant distortion when $q\leq\frac{k}{2}$, implying that this known impossibility from the setting with disjoint voters and candidates and a general metric space (Caragiannis et al., 2022b) remains in place in our restricted setting. To prove this bound, we partition all but $q$ agents into $\big\lfloor\frac{k}{q}\big\rfloor\geq 2$ sets and consider two metrics that differ in the position of the remaining $q$ agents: relatively close to the other agents in one metric; very far in the other. Intuitively, selecting these $q$ agents leads to an unbounded distortion in the former case but is necessary for a bounded distortion in the latter. For $q>\frac{k}{2}$, the existence of rules with distortion $3$ follows from a general result by Caragiannis et al. (2022b). We provide a lower bound that varies between $\frac{3}{2}$ and $2$ as $q$ varies between $\frac{k}{2}+1$ and $2$, by considering three different metrics consistent with the same rankings and showing that, in one of them, there are $q$ agents in one extreme that cannot be consistently selected. We take a closer look at the case with $k=q=2$, where a best-possible distortion of $2$ can be achieved by selecting the median agents when $n$ is even. For odd $n$, we show that a rule selecting a couple of agents—a pair of agents who prefer each other over all other agents—among the five middle agents achieves an improved distortion of $\frac{4}{3}$, which is again best-possible. The Favorite Couple rule leverages two key principles: (1) selecting agents close to the median so as to balance overall distances from agents on each side of the median, and (2) selecting consecutive agents with a small distance between them since this distance is part of the cost of all agents. This intuition of selecting consecutive agents that are as close to each other as possible while also being close to the median in principle holds for larger $k$, but determining how tightly a group of $k$ agents is clustered based solely on ordinal rankings remains a challenge.
@@
In Section 4.3, we show that $k$-Extremes attains a distortion of 22 for $q$-cost as long as $q>\frac{k}{3}$. To do so, we prove that the social cost of the set selected by this rule is at most the distance from the agent closest to the center to their nearest extreme, and bound the social cost of the optimal set from below by half of this distance. We provide a matching lower bound by revisiting the instance used for the additive case. Finally, we show that no constant distortion is possible when $q\leq\frac{k}{3}$, again implying that the general impossibility result of Caragiannis et al. (2022b) still holds in our setting. In the worst-case instances, we partition the agents into $\big\lfloor\frac{k}{q}\big\rfloor$ sets and consider two symmetric distance metrics where all but one set are placed at a unit distance from one another and two sets in one extreme are at the same location. We show that no rule can pick $q$ agents from each location.
@@ 1.2 Further Related Work
In the non-metric framework, Caragiannis and Procaccia (2011) showed that the distortion of any voting rule is at least $\Omega(m^{2})$ and that simple rules such as Plurality achieve a distortion of at most $O(m^{2})$, where $m$ is the number of candidates. In the metric framework, Anshelevich et al. (2018) established a general lower bound of $3$ on the distortion of any deterministic voting rule. They also analyzed the distortion of common voting rules such as Majority, Borda, and Copeland, showing that the latter achieves the lowest distortion of $5$ among them. Goel et al. (2017) disproved a conjecture by Anshelevich et al. regarding a better-than-$5$ distortion of the Ranked Pairs rule and introduced the notion of fairness ratio of a rule, which captures the egalitarian social cost as a special case. These results were later improved by Munagala and Wang (2019), who extended the analysis to uncovered set rules and reduced the upper bound to $4.236$. Gkatzelis et al. (2020) closed the gap by improving this bound to $3$, and showed the validity of this bound in terms of fairness ratio and thus egalitarian social cost. Randomized voting rules have also been extensively explored in the metric framework (Pulyassary and Swamy, 2021; Fain et al., 2019). The best-known upper bound for a randomized voting rule was recently obtained by Charikar et al. (2024), who showed that a carefully designed randomization over existing and novel voting rules achieves a distortion of at most $2.753$. As for lower bounds, Charikar and Ramakrishnan (2022) disproved a conjecture by Goel et al. (2017) regarding the existence of a randomized voting rule with distortion $2$, by constructing instances whose distortion approaches $2.113$ as the number of candidates grows.
@@
In the study of metric distortion for multi-winner voting, various objective functions have been proposed to capture the cost incurred by each voter for the elected committee (Elkind et al., 2017; Faliszewski et al., 2017). A foundational result by Goel et al. (2018) showed that, for the additive cost function, iterating a single-winner voting rule with distortion $\delta$ for $k$ rounds produces a $k$-winner committee with the same distortion. Chen et al. (2020) studied the $1$-cost objective in the metric framework when each voter casts a vote for a single candidate. They proposed a deterministic rule with a tight distortion of $3$ and a randomized rule with a distortion of $3-\frac{2}{m}$. More generally, Caragiannis et al. (2022b) introduced the $q$-cost objective, where a voter’s cost for a committee is determined by the distance to their $q$th closest member. They showed that the distortion is unbounded for $q\leq\frac{k}{3}$ and linear in $n$ for $\frac{k}{3}<q\leq\frac{k}{2}$. For $q>\frac{k}{2}$, they presented a non-polynomial voting rule that achieves a distortion of $3$ and a polynomial rule with a distortion of $9$. They discussed how these upper bounds for $q>\frac{k}{2}$ and the unbounded distortion for $q\leq\frac{k}{3}$ carry over to egalitarian social cost, but interestingly showed that a constant distortion is possible for this objective when $\frac{k}{3}<q\leq\frac{k}{2}$. Kizilkaya and Kempe (2022) later proposed a polynomial-time rule with a distortion of $3$. Recently, Babashah et al. studied the distortion of multi-winner elections with additive cost on the line, devising a rule with a distortion of roughly $\frac{7}{3}$. Caragiannis et al. (2017) studied distortion in multi-winner voting for the non-metric framework, defining a voter’s utility for a committee as the highest utility derived from any of its members. They proposed a rule achieving a distortion of $1+\frac{m(m-k)}{k}$ for deterministic committee selection when selecting $k$ out of $m$ candidates.
@@
A specialized setting in metric voting studies single-peaked and $1$-Euclidean preferences, where both voters and alternatives are embedded on the real line (Black, 1948; Moulin, 1980; Miyagawa, 2001; Fotakis et al., 2016; Fotakis and Gourvès, 2022; Voudouris, 2023; Ghodsi et al., 2019). In particular, the work of Fotakis et al. (2024) investigated the distortion of deterministic algorithms for $k$-committee selection on the line under the $1$-cost objective, leveraging additional distance queries.